f(x)=cotx+tanx²-2cosx²


1
Expert's answer
2020-02-04T08:51:09-0500

"\frac{df}{dx}=-\frac{1}{sin^2(x)}+\frac{2x}{cos^2(x^2)}-2\cdot (-sin(x^2)\cdot 2x)="

"=-\frac{1}{sin^2(x)}+\frac{2x}{cos^2(x^2)}+4xsin(x^2)"